OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 4 experience points. On level 3, a player must earn 8 experience points. On level 4, a player must earn 13 experience points. How many experience points must a player earn on level 7?

OpenStudy (anonymous):

I dont know how to do this :/

OpenStudy (anonymous):

from 1 to 4 add 3 from 4 to 8 add 4 from 8 to 13 add 5

OpenStudy (anonymous):

next one would be \(13+6=19\)

OpenStudy (anonymous):

then \(19+7=26\)

OpenStudy (anonymous):

and finally \(26+8=34\)

OpenStudy (anonymous):

well... that was really easy... i forgot to think a lil bit :/

OpenStudy (anonymous):

there is probably a fancier way to do it, as this is a quadratic function of some kind

OpenStudy (anonymous):

yah... i was stuck by looking for some ecuation tho.. but i could have done it as eas as u did.. thanks bro

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

actually the equation is not that easy to find it is \[\frac{x^2}{2}+\frac{x}{2}-1\]